### v2

# Validation tests of Gaussian boson samplers with photon-number resolving detectors

Abstract
An important challenge with the current generation of noisy, large-scale quantum computers is the question of validation. Does the hardware generate correct answers? If not, what are the errors? This issue is often combined with questions of computational advantage, but it is a fundamentally distinct issue. In current experiments, complete validation of the output statistics is generally not possible because it is exponentially hard to do so. Here, we apply phase-space simulation methods to partially verify recent experiments on Gaussian boson sampling (GBS) implementing photon-number resolving (PNR) detectors. The positive-P phase-space distribution is employed, as it uses probabilistic sampling to reduce complexity. It is $\sim 10^{18}$ times faster than direct classical simulation for experiments on $288$ modes where quantum computational advantage is claimed. When combined with binning and marginalization to improve statistics, multiple validation tests are efficiently computable, of which some tests can be carried out on experimental data. We show that the data as a whole shows discrepancies with theoretical predictions for perfect squeezing. However, a small modification of the GBS parameters greatly improves agreement for some tests. We suggest that such validation tests could form the basis of feedback methods to improve GBS experiments.
I Introduction
Theoretical and experimental interest in photonic networks employed as quantum computers has grown considerably [1, 2]. These devices interfere photons from nonclassical states in a large scale interferometric network, then measure the resulting output photons using photodetectors. Of the proposed network architectures, Gaussian boson sampling (GBS) has proved scalable to the sizes needed to infer quantum advantage, which recent experiments have now claimed [3, 4, 5, 6]. In this case, photons are prepared in a squeezed state and the computational task is sampling from an output distribution with probabilities in a #P-hard computational class [7, 8]. For photon-number resolving (PNR) detectors this is a Hafnian, while for threshold detectors it is a Torontonian. Neither function is computable in polynomial time, so that the random counts cannot be classically replicated at large scale.
A crucial question with all current generation quantum computers is how to test for errors in the computed outputs. An efficient and scalable method for validating GBS with threshold detectors, demonstrated for up to $16,000$ modes, was proposed in Ref.[9] and used to analyze data from a $100$-mode experiment [3]. The normally ordered positive-P phase-space representation was utilized to simulate grouped count probabilities (GCPs). This is a generalized binned photon counting distribution that includes all measurable marginal and binned correlations. The exponential sparsity of the detected count pattern probabilities means the full distribution cannot be either computed or measured. Therefore, validation tests based on marginals and binned probability distributions are the most practical method for validation, as is the case for other random number generators [10].
In this paper, we use positive P-representation simulations to test the validity of Gaussian boson samplers with photon number resolving (PNR) detectors, extending the tests that can be applied to experimental data. This is essential, as a number of large scale GBS experiments have claimed quantum advantage using either PNR [5] or threshold detectors [3, 4, 6]. To perform the tests, we generalize a binning method developed by Mandel [11, 12] for simulating photon statistics of classical light. The original formulation, which utilized the Glauber-Sudarshan P-representation [13, 14] to simulate moments for classical fields, is extended to the positive P-representation, which can treat any quantum state. An extension to binning in multiple dimensions is also provided, as these are more sensitive tests than one-dimensional distributions.
We show that in the challenging prospect of using quantum computers to demonstrate quantum computational advantage, the GBS approach has a very important advantage. Unlike many other proposed quantum computing architectures, there are multiple quantitative statistical tests that are dependent on the experimental parameters, which can be used to validate the output distributions of the data. Experimental data may fail such validation due to errors, potentially allowing classical simulations of comparable accuracy. This is an important criterion, similar to tests of classical random number generators and cryptographic security [15, 16]. Implementing a test suite requires algorithms that compute the correlations or moments without requiring the full #P hard distribution.
Like phase-space methods for threshold detectors [9], the PNR method is efficient and scalable. Increased dimension of binning causes an increase in computation time, where in the limit of dimension equaling mode number the GCP converges to the Hafnian. PNR phase-space methods allow one to perform simulations of either the ground truth distribution or modifications of this distribution that include decoherence, measurement errors and shot-to-shot fluctuations. Such errors are typically not analyzed, despite evidence suggesting that a thermalized, non-ideal ground truth distribution provides stronger claims of quantum advantage than the ideal ground truth distribution [17]. Unlike threshold detection, PNR detector binning has the advantage of not requiring the computation of multidimensional inverse discrete Fourier transforms, which can increase computation time for increased dimension.
The positive-P phase-space distribution is exact, non-singular and positive for any quantum state. The moments of the distribution are identical to normally ordered operator moments including intensity correlations [18], which are reproduced to any order. This property, coupled with probabilistic random sampling, makes the positive P-representation ideal for validating the statistics of nonclassical state photon counting experiments such as GBS. It does not have the computational restraints imposed by other validation methods [19, 20, 21], since it is many orders of magnitude faster than direct classical simulation, with negligible rounding errors. While there are sampling errors, for current experimental parameters they are comparable to sampling errors in the measurement statistics. Both can be reduced in a controllable way.
Other validation methods involve comparing experimental count patterns to those obtained from classical algorithms that replicate the photo-detection measurements either exactly [21] or approximately [19, 20]. Exact algorithms generate count patterns by directly sampling the distributions, but are limited to small experiments only. Approximate algorithms typically neglect high-order correlations, and only use the low-order correlations. Despite the usefulness of these algorithms in defining classical simulation boundaries of quantum advantage and determining the correlations present in a GBS experiment, they are either computationally demanding or lack high-order correlations, which is necessary for claims of quantum advantage. This limits the application of these approaches when validating GBS experiments and investigating physical processes arising within networks.
In order to determine the validity of using the positive P-representation to simulate photon counting experiments with PNR detectors, we compare simulated GCP moments to the exactly known photon counting distributions for special cases. Here, multi-mode pure squeezed states are transformed by a Haar random unitary matrix which is multiplied by a uniform amplitude loss coefficient. For losses comparable to current experiments, we see excellent agreement. In the less realistic lossless case, the resulting non-Gaussian number distribution oscillates strongly between even and odd total counts. While this is far from current experiments, the standard positive P-representation requires a much larger number of phase-space samples to converge in such cases. However, this limit can also be efficiently simulated if necessary using a symmetry-projected [22] phase-space method, which will be treated elsewhere.
Finally, using photon count data from the recent Borealis experiment implementing PNR detectors [5], we apply the simulated GCPs to validate the moments of measured photon counting distributions by comparing theory and experiment. Differences are quantified using chi-square and Z-statistic tests. As with earlier work [9, 17], we show that the experimental GCPs are far from the ground truth distribution for pure squeezed state inputs. Instead, they are described much better by a different target distribution which includes decoherence and measurement error corrections of the transmission matrix. Although we focus specifically on implementations to GBS, we emphasize that one can use the positive P-representation to simulate GCPs of any linear photon counting experiment using classical or nonclassical state photons.
This paper is structured as follows, Sections II and III introduce the necessary background and theoretical methods for simulating GCPs of any photon counting experiment using PNR detectors in phase-space. Section IV then applies this method to the exactly known multi-mode squeezed state distributions, while Section V simulates the theoretical distributions corresponding to experimental data from the Borealis GBS claiming quantum advantage. Finally, VI summarizes the main findings of this paper.
II Grouped count probabilities
We first consider a standard GBS experiment where $N$ single-mode squeezed vacuum states are sent into a linear photonic network defined by an $M\times M$ transmission matrix $\boldsymbol{T}$. If each input mode is independent, the full input density operator is defined as
$\hat{\rho}^{(\text{in})}=\prod_{j}^{N}\left|r_{j}\right\rangle\left\langle r_{j}\right|,$ (1)
where $\left|r_{j}\right\rangle=\hat{S}(r_{j})\left|0\right\rangle$ denotes the squeezed vacuum state of the $j$-th mode with squeezing parameter $r_{j}$ and $\hat{S}(r_{j})$ is the squeezing operator [23].
Since squeezed states are Gaussian, they are characterized by their quadrature variances
$\left\langle:\left(\Delta\hat{x}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}+m_{j}\right)=e^{2r_{j}}-1$
$\left\langle:\left(\Delta\hat{y}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}-m_{j}\right)=e^{-2r_{j}}-1,$ (2)
where the quadrature operators $\hat{x}_{j}=\hat{a}_{j}^{(\text{in})}+\hat{a}_{j}^{\dagger(\text{in})}$ and $\hat{y}_{j}=\left(\hat{a}_{j}^{(\text{in})}-\hat{a}_{j}^{\dagger(\text{in})}\right)/i$ satisfy the commutation relations $\left[\hat{x}_{j},\hat{y}_{k}\right]=2i\delta_{jk}$, and the mean photon number and coherence per mode is defined as $n_{j}=\left\langle\hat{a}_{j}^{\dagger(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\sinh^{2}(r_{j})$ and $m_{j}=\left\langle\hat{a}_{j}^{(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\cosh(r_{j})\sinh(r_{j})$, respectively. Throughout this paper, we use ::$\cdots$:: ⋯ : to denote normal ordering.
While most theoretical analyses of GBS consider pure squeezed state inputs, this is impractical, since decoherence effects such as mode mismatches [3], partial distinguishability [24], and many more [25] will always be present during experimental applications in quantum optics. Therefore, a realistic model of input states are Gaussian thermalized squeezed states for which a simple parameterization was introduced in [9]. Here, a thermalization component $\epsilon$ alters the intensity of the input photons as measured by the coherence, which is reduced by a factor of $1-\epsilon$ such that $\tilde{m}_{j}=(1-\epsilon)m_{j}$. The quadrature variances are altered as [9]
$\sigma_{x_{j}}^{2}=\left\langle:\left(\Delta\hat{x}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}+\tilde{m}_{j}\right)$
$\sigma_{y_{j}}^{2}=\left\langle:\left(\Delta\hat{y}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}-\tilde{m}_{j}\right),$ (3)
where for $\epsilon=0$, the pure squeezed state definition Eq.(2) returns, while $\epsilon=1$ produces fully decoherent thermal states with $\sigma_{x_{j}}^{2}=\sigma_{y_{j}}^{2}=2n_{j}$. More complex thermalized models are also possible, for example with mode dependent thermalization, $\epsilon_{j}$.
In the ideal GBS scenario, the transmission matrix is a Haar random unitary. However experimentally realistic transmission matrices contain losses, making them non-unitary. Throughout this paper, $\boldsymbol{T}$ is used to define a lossy transmission matrix, while $\boldsymbol{U}$ defines a Haar random unitary. When losses are present, the creation and annihilation operators for the input modes $\hat{a}_{j}^{\dagger(\text{in})},\hat{a}_{j}^{(\text{in})}$ are converted to output modes following
$\hat{a}_{k}^{\dagger(\text{out})}=\sum_{j=1}^{N}T_{kj}\hat{a}_{j}^{\dagger(\text{in})}+\sum_{j=1}^{M}L_{kj}\hat{b}_{j}^{\dagger(\text{in})},$ (4)
where $\boldsymbol{L}$ is a random noise matrix representing photon loss to $j$-the environmental mode $\hat{b}_{j}^{\dagger(\text{in})}$. The resulting output photon number $\hat{n}^{\prime}_{k}=\hat{a}_{k}^{\dagger(\text{out})}\hat{a}_{k}^{(\text{out})}$ is then measured by a series of photodetectors.
II.1 Photon counting projectors
Detectors with photon number resolution implement the normally ordered projection operator [26, 27]
$\hat{p}_{j}(c_{j})=\frac{1}{c_{j}!}:\left(\hat{n}^{\prime}_{j}\right)^{c_{j}}e^{-\hat{n}^{\prime}_{j}}:,$ (5)
where $c_{j}=0,1,2,3,\dots$ is the number of measured photon counts at the $j$-th detector. Although theoretically there is no upper bound on $c_{j}$, practically PNR detectors saturate for some maximum number of counts $c_{j}^{(\text{max})}$. This value is variable, typically depending on the physical implementation of the detector.
In any multi-mode photon counting experiment, a set of photon count patterns $\boldsymbol{c}=[c_{1},\dots,c_{M}]$ are observed by measuring the output state $\hat{\rho}^{(\text{out})}$. A straightforward extension of Eq.(5) allows one to define the projection operator for a specific count pattern as
$\hat{P}(\boldsymbol{c})=\bigotimes_{j=1}^{M}\hat{p}_{j}(c_{j}).$ (6)
For GBS, each pattern is a sample from the full output probability distribution and computing its expectation value corresponds to solving the matrix Hafnian [7, 28]
$\left\langle\hat{P}(\boldsymbol{c})\right\rangle=\frac{1}{\sqrt{\text{det}(\boldsymbol{Q})}}\frac{\left|\text{Haf}\left(\boldsymbol{B}_{S}\right)\right|^{2}}{\prod_{j=1}c_{j}!},$ (7)
which is a #P-hard computational task due to its relationship to the matrix permanent. Here, $\text{Haf}(\boldsymbol{B}_{S})$ is the Hafnian of the square sub-matrix of $\boldsymbol{B}=\boldsymbol{U}\left(\bigoplus_{j=1}^{M}\tanh(r_{j})\right)\boldsymbol{U}^{T}$ for pure state inputs with zero displacement, where the sub-matrix is formed from the output channels with recorded photon counts. The symmetric matrix $\boldsymbol{B}$ is related to the kernel matrix $\boldsymbol{A}=\boldsymbol{B}\bigoplus\boldsymbol{B}^{*}$ which is formed from the $2M\times 2M$ covariance matrix $\boldsymbol{Q}$ as [29]
$\boldsymbol{A}=\left(\begin{array}[]{cc}0&\boldsymbol{I}_{M}\\ \boldsymbol{I}_{M}&0\end{array}\right)\left(\boldsymbol{I}_{M}-\boldsymbol{Q}^{-1}\right),$ (8)
where $\boldsymbol{I}_{M}$ denotes the $M\times M$ identity matrix. When the covariance matrix elements are non-negative, the kernel matrix and hence the symmetric matrix $\boldsymbol{B}$ are also non-negative [29] and the Hafnian can be estimated in polynomial time [30, 31].
In the current generation of GBS networks, photon loss is the dominant noise source. Although there is evidence that sampling from a lossy GBS distribution is also a #P-hard computational task [32], output photons from experiments implementing high loss networks become classical [33], allowing the output distribution to become efficiently sampled. There is currently no clear boundary between the computationally hard and efficiently sampled regimes, but a classical algorithm that exploits large experimental photon loss has recently been developed [34] and shown to produce samples that are closer to the ideal ground truth than recent experiments claiming quantum advantage. The effects of thermalization on the computational complexity of GBS is currently unknown, despite thermalization causing photons from nonclassical states to become classical at a faster rate than pure losses.
II.2 Grouped count probabilities for PNR detectors
Determining photon counting distributions by binning the photons arriving at photodetectors has a long history in quantum optics. Although binning methods can vary, e.g. binning continuous measurements within a time interval such as inverse bandwidth time [35, 36], such distributions are both experimentally accessible and theoretically useful. For pure squeezed states these can provide evidence of nonclassical behavior, as observed in the even-odd count oscillations [37]. For GBS there is also a statistical purpose, as the set of all possible observable patterns $\mathcal{S}_{P}$ grows exponentially with mode number, causing the output probabilities to become exponentially sparse.
In this paper, we are interested in the total output photon number operator obtained from a subset of $S_{j}$ efficient detectors:
$\hat{n}^{\prime}_{S_{j}}=\sum_{i\in S_{j}}\hat{n}^{\prime}_{i}.$ (9)
Our observable of interest is then the photon number distribution corresponding to counting the projectors onto eigenstates of $\hat{n}^{\prime}_{S_{j}}$.
Following from the nomenclature introduced in [9], we define the $d$-dimensional grouped photon counting distribution as the probability of observing $\boldsymbol{m}=[m_{1},\dots,m_{d}]$ grouped counts in $j=1,\dots,d$ distinct subsets of output modes $\boldsymbol{S}=\text{$\left(S_{1},\dots,S_{d}\right)$}$, such that
$\mathcal{G}_{\boldsymbol{S}}^{(n)}(\boldsymbol{m})=\left\langle\prod_{j=1}^{d}\left[\sum_{\sum c_{i}=m_{j}}\hat{P}_{S_{j}}(\boldsymbol{c})\right]\right\rangle,$ (10)
where $n=\sum_{j=1}^{d}M_{j}\leq M$ is the total correlation order and each grouped count is obtained from the summation $m_{j}=\sum_{i\in S_{j}}c_{i}$. The general form of the pattern projector for any number of output modes within each subset is
$\hat{P}_{S_{j}}\left(\boldsymbol{c}\right)=\bigotimes_{i\in S_{j}}\frac{1}{c_{i}!}:(\hat{n}^{\prime}_{i})^{c_{i}}e^{-\hat{n}^{\prime}_{i}}:.$ (11)
The correlation order and dimension determines the type of observable one is simulating. If $n=M$, all correlation orders are simulated in phase-space. Meanwhile $n<M$ denotes a marginal probability, since the remaining $M-n$ output modes are ignored. If $n=M$ and $d=1$, the GCP becomes the univariate distribution $\mathcal{G}_{S}^{(M)}(m)$ where all modes are contained within a single subset $S=\{1,\dots,M\}$. This is called the total count distribution as it’s the probability of observing $m$ counts in any pattern and is identical to the total photon-number distribution of Ref. [5].
When $d>1$, multi-dimensional GCPs are simulated. These are joint probability distributions over $\boldsymbol{m}$ counts. For example, one of the $d=2$ GCPs is the bivariate photon count distribution $\mathcal{G}_{S_{1},S_{2}}^{(M)}(m_{1},m_{2})$, which is the probability of observing $m_{1}$ counts in $S_{1}=\{1,\dots,M_{1}\}$, and $m_{2}$ counts in $S_{2}=\{M_{1}+1,\dots,M\}$. The largest possible dimension is $d=M$ where the vector of subsets becomes $\boldsymbol{S}=(S_{1},S_{2},\dots,S_{M})=(\{1\},\{2\},\dots,\{M\})$. In this case, the GCP to converges to the Hafnian as the joint probabilities have become pattern probabilities.
Provided one has the computational resources (see Section IV.3), comparisons of multi-dimensional GCPs provide a number of benefits over their $d=1$ counterparts. The first is the number of count bins scales as $\mathcal{B}=(Md^{-1}+1)^{d}$, allowing more detailed comparisons with theory and experiment. The second is by randomly permuting the measured count patterns, an exponentially large number of comparison tests can be performed on experimental data, where each permutation allows one to compare different binned correlation moments [17].
So long as $n=M$ all correlation orders are simulated in phase-space. However, since GCP probabilities converge to the Hafnian as $d\rightarrow M$, a greater range of distinct correlations are statistically significant in this limit. If these correlations are nonclassical, multi-dimensional GCPs provide a more thorough test of quantum advantage, because they allow one to directly compare more sums of high-order correlations with their theoretical values [17], and these tests are more sensitive to the exact squeezing and transfer matrix parameter values.
III Normally-ordered phase-space methods
Here we summarize phase-space methods, and explain the motivation for using them. The purpose is to transform an exponentially complex Hilbert space into a positive distribution that can be randomly sampled. Normal-ordering methods are the closest to experimental statistics using photon-counting. Representations that are not normally ordered, like the Wigner [38] or Q-function [39] approach, have a rapidly growing sampling error with system size.
III.1 Glauber-Sudarshan representation
In its original formulation, Mandel [11, 12] used the normally ordered Glauber-Sudarshan P-representation [13, 14] to evaluate the expectation value in Eq.(10). The P-representation expands the density operator as a diagonal sum of coherent states [13]
$\hat{\rho}=\int P(\boldsymbol{\alpha})\left|\boldsymbol{\alpha}\right\rangle\left\langle\boldsymbol{\alpha}\right|\text{d}^{2M}\boldsymbol{\alpha},$ (12)
hence its commonly called the diagonal P-representation, where $\left|\boldsymbol{\alpha}\right\rangle=\bigotimes_{j=1}^{M}\left|\alpha_{j}\right\rangle$ is a multi-mode coherent state vector with eigenvalues $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{M}]^{T}$ and $P(\boldsymbol{\alpha})$ is the P-representation probability distribution. The diagonal P-representation can be used to evaluate expectation values of the product of normally ordered operators due to the relationship
$\left\langle\prod_{j=1}^{M}:\hat{a}_{j}^{\dagger}\hat{a}_{j}:\right\rangle$ $=\lim_{E_{S}\rightarrow\infty}\left\langle\prod_{j=1}^{M}\left|\alpha_{j}\right|^{2}\right\rangle_{DP}$
$=\int P(\boldsymbol{\alpha})\left(\prod_{j=1}^{M}\left|\alpha_{j}\right|^{2}\right)\text{d}^{2M}\boldsymbol{\alpha},$ (13)
[…]
Following the standard replacement of $\hat{a}_{j}\rightarrow\alpha_{j}$ and $\hat{a}_{j}^{\dagger}\rightarrow\alpha_{j}^{*}$, where $n^{\prime}_{j}=\left|\alpha_{j}\right|^{2}$ and $\hat{n}^{\prime}_{S_{j}}\rightarrow n^{\prime}_{S_{j}}=\sum_{i\in S_{j}}\left|\alpha_{i}\right|^{2}$, Mandel [11, 12] simulates Eq.(10) for $d=1$ and $n=M$ as
$\mathcal{G}_{S}^{(M)}(m)=\int P(\boldsymbol{\alpha})\left[\frac{1}{m!}\left(n^{\prime}_{S}\right)^{m}e^{-n^{\prime}_{S}}\right]\text{d}^{2M}\boldsymbol{\alpha}.$ (14)
Since the density operator of nonclassical states contains off-diagonal elements due to quantum superpositions, if the diagonal P-representation is used for simulations of nonclassical states, $P(\boldsymbol{\alpha})$ becomes singular and negative. From standard mathematical requirements that it should be real, normalizable, non-negative and non-singular, it is no longer a probability distribution. This method is therefore only useful for classical states such as coherent, thermal and squashed states [40, 41], since these have a positive diagonal P-distribution.
III.2 Positive-P representation
To simulate the photon number distributions of nonclassical states one must instead use the normally ordered positive P-representation [42], which is part of a family of generalized P-representations that produce non-singular and exact phase-space distributions for any quantum state. The positive P-representation expands the density operator as a $4M$-dimensional volume integral [42]
$\hat{\rho}=\iint P(\boldsymbol{\alpha},\boldsymbol{\beta})\frac{\left|\boldsymbol{\alpha}\right\rangle\left\langle\boldsymbol{\beta}\right|}{\left\langle\boldsymbol{\beta}|\boldsymbol{\alpha}\right\rangle}\text{d}^{2M}\boldsymbol{\alpha}\text{d}^{2M}\boldsymbol{\beta},$ (15)
where the additional dimensions allow off-diagonal density matrix elements to be included in the coherent state basis $\left|\boldsymbol{\beta}\right\rangle=\bigotimes_{j=1}^{M}\left|\beta_{j}\right\rangle$ with eigenvalues $\boldsymbol{\beta}=[\beta_{1},\dots,\beta_{M}]^{T}$. The coherent amplitudes $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$ are independent and can vary along the entire complex plane. We can also take the real part of Eq.(15), since the density matrix is hermitian. The benefit of using the positive P-representation for photon counting experiments is that it includes the diagonal P-representation when the positive-P distribution satisfies $P(\boldsymbol{\alpha},\boldsymbol{\beta})=P(\boldsymbol{\alpha})\delta(\text{$\boldsymbol{\alpha}$-$\boldsymbol{\beta}$})$. This allows us to define classical phase-space as having $\boldsymbol{\beta}=\boldsymbol{\alpha}$, while a nonclassical phase space requires non-vanishing probability of $\boldsymbol{\beta}\neq\boldsymbol{\alpha}$.
Expectation values of products of normally ordered operators for nonclassical states can now be directly calculated from the moments of the positive-P distribution
$\left\langle\prod_{j=1}^{M}:\hat{a}_{j}^{\dagger}\hat{a}_{j}:\right\rangle$ $=\lim_{E_{S}\rightarrow\infty}\left\langle\prod_{j=1}^{M}\alpha_{j}\beta_{j}^{*}\right\rangle_{P}$
$=\iint P(\boldsymbol{\alpha},\boldsymbol{\beta})\left(\prod_{j=1}^{M}\alpha_{j}\beta_{j}^{*}\right)\text{d}^{2M}\boldsymbol{\alpha}\text{d}^{2M}\boldsymbol{\beta},$ (16)
where we define $\left\langle\dots\right\rangle_{P}$ as the positive-P ensemble mean and have used the c-number replacements $\hat{a}_{j}\rightarrow\alpha_{j}$ and $\hat{a}_{j}^{\dagger}\rightarrow\beta_{j}^{*}$.
To simulate Eq.(10) for nonclassical states requires replacing the total output photon number Eq.(9) by the phase-space observable
$n^{\prime}_{S_{j}}=\sum_{i\in S_{j}}n^{\prime}_{i}=\sum_{i\in S_{j}}\alpha^{\prime}_{i}\left(\beta^{\prime}_{i}\right)^{*},$ (17)
such that the general multidimensional probability distribution is
$\mathcal{G}_{\boldsymbol{S}}^{(n)}(\boldsymbol{m})$ $=\left\langle\prod_{j=1}^{d}\left[\sum_{\sum c_{i}=m_{j}}P_{S_{j}}(\boldsymbol{c})\right]\right\rangle_{P},$
$=\text{Re}\int P(\boldsymbol{\alpha},\boldsymbol{\beta})\left[\prod_{j=1}^{d}\left[\frac{1}{m_{j}!}\left(n^{\prime}_{S_{j}}\right)^{m_{j}}e^{-n^{\prime}_{S_{j}}}\right]\right]\text{d}\mu,$ (18)
where $\text{d}\mu\equiv\text{d}\boldsymbol{\alpha}\text{d}\boldsymbol{\beta}$ and the phase-space pattern projector is
$P_{S_{j}}(\boldsymbol{c})=\prod_{i\in S_{j}}\frac{1}{c_{i}!}(n^{\prime}_{i})^{c_{i}}e^{-n^{\prime}_{i}}.$ (19)
In Eq.(18), the required summation over output photon number is incorporated into the pattern projector, allowing one to make the substitution $m_{j}=\sum_{i\in S_{j}}c_{i}$, while we take the real part as required for pure squeezed state applications presented in the next subsection.
Equation 18 is a more general definition of GCPs than Mandel’s original formulation, since one can now simulate the marginal and multidimensional probabilities of any classical or nonclassical photon counting experiment. To simulate GCPs for input photons in a number state, it is more efficient to use another generalized P-representation: the complex P-representation. In this case $P(\boldsymbol{\alpha},\boldsymbol{\beta})$ includes a complex weight due to expanding the density operator as a contour integral [42]. Although the positive P-representation exists for Fock states, it has larger sampling errors than the complex version [43], meaning fewer stochastic samples are needed to perform accurate simulations.
Unlike methods for threshold detectors [9], Eq.(18) doesn’t require multidimensional inverse discrete Fourier transforms, which can increase the computation time. Since threshold detector outputs are always binary, the largest number of observable photon counts in a pattern corresponds to a detection event in every output mode, such that $\sum_{i}^{M}c_{i}\leq M$ where the largest observable grouped count is $m^{(M)}$.
Since PNR detectors can resolve multiple photons arriving at each detector, an experiment or classical algorithm replicating photo-detection measurements can now produce count patterns where $\sum_{i}^{M}c_{i}>M$, in which case $m^{(M)}$ is no longer the largest observable grouped count. Instead of scaling with mode number, numerical simulations scale with photon count bins, where we define the maximum observable grouped count as $m^{(\text{max})}$. For experiments with small mean input photon numbers or high losses, one may have $m^{(\text{max})}<m^{(M)}$, while low losses and large photon numbers will produce $m^{(\text{max})}>m^{(M)}$.
Typically, we choose $m^{(\text{max})}$ to occur when probabilities are of the order $\mathcal{G}_{S}^{(n)}(m^{(\text{max})})\leq 10^{-7}$, providing a standard cutoff for very small probabilities as was implemented in previous methods [9, 17]. A minimum observable grouped count $m^{(\text{min})}$ can also be introduced to form an analogous cut-off $\mathcal{G}_{S}^{(n)}(m^{(\text{min})})\leq 10^{-7}$. This allows our algorithm to only simulate grouped counts within the relevant range $m_{j}^{(\text{min})}\leq m_{j}\leq m_{j}^{(\text{max})}$, reducing memory use and computation times.
III.3 Squeezed states in phase-space
Initial coherent amplitudes are generated by randomly sampling the phase-space distribution of a specific state. For most states, analytical forms of this distribution exist, allowing one to derive an analytical expression for the initial samples of each state. Here we simply recall previously derived results for squeezed states [9], so that the theory presented is complete, since the form of the positive-P distribution doesn’t depend on the type of measurement, that is, PNR or not.
For applications to GBS, we first consider the ideal case with pure squeezed state inputs. In order to define the input density operator Eq.(1) in the positive P-representation we first expand each state as an integral over real coherent states [44]
$\left|r_{j}\right\rangle=\sqrt{C_{j}}\int\exp\left(\frac{-\alpha_{j}^{2}}{\gamma_{j}}\right)\left|\alpha_{j}\right\rangle\text{d}\alpha_{j},$ (20)
with normalization constant $C_{j}=\sqrt{\gamma_{j}+1}/(\pi\gamma_{j})$ and $\gamma_{j}=\exp(2r_{j})-1$. After some straightforward calculations, the input density operator becomes
$\hat{\rho}^{(\text{in})}=\text{Re}\iint P(\boldsymbol{\alpha},\boldsymbol{\beta})\frac{\left|\boldsymbol{\alpha}\right\rangle\left\langle\boldsymbol{\beta}\right|}{\left\langle\boldsymbol{\beta}|\boldsymbol{\alpha}\right\rangle}\text{d}\boldsymbol{\alpha}\text{d}\boldsymbol{\beta},$ (21)
where the positive-P distribution for pure squeezed states is [9]
$P(\boldsymbol{\alpha},\boldsymbol{\beta})=\prod_{j}C_{j}e^{-\left(\alpha_{j}^{2}+\beta_{j}^{2}\right)\left(\gamma^{-1}+1/2\right)+\alpha_{j}\beta_{j}}.$ (22)
Thermalized squeezed states, defined by the $\epsilon$ component introduced in Section II, are readily incorporated into the positive-P distribution. A general random sampling formalism is used to generate initial stochastic samples for any Gaussian input [9]
$\alpha_{j}$ $=\frac{1}{2}\left(\sigma_{x_{j}}w_{j}+i\sigma_{y_{j}}w_{j+M}\right)$
$\beta_{j}^{*}$ $=\frac{1}{2}\left(\sigma_{x_{j}}w_{j}-i\sigma_{y_{j}}w_{j+M}\right),$ (23)
where $\left\langle w_{j}w_{k}\right\rangle_{P}=\delta_{jk}$ are real Gaussian noises and the variances $\sigma_{x_{j}}$, $\sigma_{y_{j}}$ are defined in Eq.(3). For classical inputs, $\sigma_{y_{j}}$ is real and $\alpha_{j}=\beta_{j}$. For quantum inputs, $\sigma_{y_{j}}$ is imaginary and $\alpha_{j}\neq\beta_{j}$. Following from Eq.(4), the output stochastic amplitudes used to perform simulations are then obtained via the transformations $\boldsymbol{\alpha}^{\prime}=\boldsymbol{T}\boldsymbol{\alpha}$ and $\boldsymbol{\beta}^{\prime}=\boldsymbol{T}\boldsymbol{\beta}$ where the amplitudes $\boldsymbol{\alpha}^{\prime}$, $\boldsymbol{\beta}^{\prime}$ correspond to measurements of the density matrix $\hat{\rho}^{\text{(out)}}$.
IV Exact models of GBS grouped count distributions
In this section, we validate the sampled positive-P method by comparing the results to exactly known distributions of GBS networks whose size, mean squeezing parameters and average loss rates correspond to those found in the PNR experiments of Madsen et al [5]. This demonstrates that for the relevant regime, the phase-space sampling errors are negligible. A full comparison with experiment that uses the inhomogeneous parameters describing the squeezing vector and transmission matrix is carried out in Section V.
The positive-P algorithm used throughout this paper is part of the phase-space network simulation program xqsim, which is available from the Github repository [45].
IV.1 Exactly known distributions
The photon statistics of pure squeezed states have been studied extensively in quantum optics [35, 36, 7, 32]. Assuming the input modes have equal squeezing parameters $r=r_{1}=\dots=r_{N}$ and hence, equal photon numbers $n=n_{1}=\dots=n_{N}$, the one-dimensional GCP, or total count probability, can be computed exactly. For the case of pure squeezed states transformed by a matrix with uniform amplitude loss coefficient $t$, applied as $t\boldsymbol{U}$, the exact total count distribution takes the form [32]
$\mathcal{G}_{S}^{(M)}(2m)$ $=t^{4m}\mathcal{C}p^{M/2}(1-p)^{m}f_{1/2}$
$\mathcal{G}_{S}^{(M)}(2m-1)$ $=2m(1-t^{2})t^{4m-2}\mathcal{C}p^{M/2}(1-p)^{m}f_{3/2},$ (24)
which is defined for $m>0$, where $S=\{1,\dots,M\}$, $p=1/(1+n)$ is the success probability, $\mathcal{C}=\binom{\frac{M}{2}+m-1}{m}$ and $f_{c}={}_{2}F_{1}\left(a,b;c;z\right)$ is the Gauss hypergeometric function with $a=m+\frac{1}{2}$, $b=\frac{M}{2}+m$, $z=(1-t^{2})^{2}(1-p)$.
For a lossless network with $t=1$, ${}_{2}F_{1}\left(a,b;c;0\right)=1$ and Eq.(24) converges to the well known distribution [35, 36, 7]
$\mathcal{G}_{S}^{(M)}(2m)$ $=\mathcal{C}p^{M/2}(1-p)^{m}$
$\mathcal{G}_{S}^{(M)}(2m-1)$ $=0,$ (25)
which contains distinct oscillations between even and odd counts. In the limit $M\rightarrow\infty$, it reduces to a Poissonian for the pair counts [36]
$\mathcal{G}_{S}^{(M)}(2m)$ $=\frac{1}{m!}e^{-Mn/2}\left(\frac{Mn}{2}\right)^{m}$
$\mathcal{G}_{S}^{(M)}(2m-1)$ $=0,$ (26)
while the full distribution is not Poissonian.
Such photon statistics are characteristic of photon bunching which, unlike photon anti-bunching, is not an exclusively nonclassical phenomena [26]. However as squeezed photons are always generated in highly correlated pairs, bunching is indicative of the non-classicality of the measured photons. The presence of even-odd oscillations [37], even for networks with decoherence and loss, provides an explicit test of non-classicality although their absence does not exclude non-classicality.
When even-odd count oscillations are present in lossless and ultra-low loss photonic networks, the positive-P moments suffer from slow convergence and large sampling errors. Since the even-odd count oscillations vanish as soon as one has a high chance of losing a photon, the lossless and ultra-low loss regimes are not relevant to any current or near future experimental network, which are both far from lossless and whose photon number distribution is approximately Gaussian for large system sizes with loss rates as small as $t=0.95$. Therefore, a detailed theoretical analysis of the positive P-representation in the lossless and ultra-low loss regimes will not be performed here, as projected methods [22] may converge faster in such regimes.
IV.2 Comparisons with small and large sized networks
To test whether Eq.(18) can be utilized to simulate nonclassical photon counting experiments using PNR detectors with perfect squeezers, photodetectors and photonic networks with uniform photon loss, we first simulate small sized GBS networks $M=N=16$ and $M=N=72$ with uniform squeezing $\boldsymbol{r}=[0.89,\dots,0.89]$ and amplitude loss rates $t=0.6$ and $t=0.56$, respectively.
Comparisons between positive-P moments and the exact distribution Eq.(24) are presented in Fig.(1a) and Fig.(1b) for these networks where simulations are performed for an ensemble size of $E_{S}=2.4\times 10^{6}$. In the limit of large $E_{S}$ the positive-P moments converge to moments of a distribution (see Eq.(16)). The exact size of $E_{S}$ required for a given accuracy varies. An ensemble size of the order $E_{S}\geq 10^{6}$ is generally needed for sampling and difference errors of less than $10^{-3}$, typical of current experimental accuracy.
For both network sizes and loss rates, the positive-P moments converge to the required exact distribution with errors $<10^{-3}$. Increasing the network sizes to $M=N=216$ and $M=N=288$, positive-P moments simulated for loss rates $t=0.57$ and $t=0.6$ are presented in Fig.(1c) and Fig.(1d). In both cases, moments converge to the exact distribution with errors $<10^{-3}$ such that they are not visible in the graphics. Therefore, in the loss regimes of the current experiments, the positive P-representation provides an accurate method to validate experimental networks. For larger sized networks within the ultra-low loss regime this is also true, as the even-odd count oscillations vanish for much smaller amplitude loss rates.
Figure 1: Validation of the positive-P GCP distributions (solid blue lines) by comparison to exact multi-mode squeezed state photon counting distributions (dashed black lines) in an exactly soluble example. We use uniform amplitude loss rates $t$ for different network sizes, squeezing parameters and loss rates corresponding to the various experimental networks implemented in Ref.[5]. a) A GBS set-up with $M=N=16$ is considered where pure squeezed states with uniform squeezing parameter $\boldsymbol{r}=[0.89,\dots,0.89]$ are transformed by a Haar random unitary matrix multiplied by a uniform amplitude loss coefficient $t\boldsymbol{U}$ with $t=0.6$. Network sizes are now increased to b) $M=N=72$, with $\boldsymbol{r}=[0.89,\dots,0.89]$ and $t=0.56$, c) $M=N=216$ with $\boldsymbol{r}=[1.1,\dots,1.1]$ and $t=0.57$, and d) $M=N=288$ with $\boldsymbol{r}=[1,\dots,1]$ and $t=0.6$. Positive-P distributions are obtained by averaging over the ensemble of samples $E_{S}=2.4\times 10^{6}$. Upper and lower lines correspond to ±$\pm 1\sigma_{T,j}$, where $\sigma_{T,j}\propto 1/\sqrt{E_{S}}$ is the theoretical sampling errors for the $j$-th photon count bin, which are so small that they are not visible in the output graphics.
IV.3 Numerical scaling
For validation tests to be useful, they must be accurate, efficient and scalable. The results above have validated the accuracy of the positive P-representation by comparing the simulated distribution to an exactly known model. Next, we analyze the efficiency and scalability of simulating GCPs using the positive P-representation for varying mode number $M$ and dimension $d$.
When performing phase-space simulations, one can subdivide the total number of samples into two sub-ensembles $E_{S}=N_{S}N_{R}$. This has both a statistical benefit, as described in the Appendix, and a computational one, as it allows both vector and multi-core parallel computation, drastically reducing computation time.
The first sub-ensemble, $N_{S}$, is the number of samples generated from a phase-space distribution, in this case Eq.(23), corresponding to a specific stochastic trajectory. The number of trajectories is denoted by second sub-ensemble $N_{R}$.
Generally, the size of each sub-ensemble affects the resulting computation time. As $N_{S}\rightarrow\infty$, large memory bandwidth to store vectors and matrices to $64$-bit (double) precision is required. Meanwhile, using $N_{R}\rightarrow\infty$ is more efficient if the hardware has more CPU cores. In practice, so long as $E_{S}$ remains unchanged, one can vary the size of each sub-ensemble to suit hardware requirements without affecting the overall accuracy of the simulation, except that the accuracy of sampling error estimates may be reduced at either extreme.
To investigate the efficiency and scalability of our method, GCPs of dimension $d=1,2,3$ are simulated for GBS networks $M=N=30,60,\dots,480$ with $\boldsymbol{r}=[1,\dots,1]$ and $t=0.6$. Positive-P numerical run-times for these networks are presented in Fig.(2) and were obtained using a typical current (2025) desktop computer with a 12-core 3.7 GHz AMD CPU and 32GB of RAM. All simulations used $E_{S}=2.4\times 10^{6}$ samples, although the sub-ensembles were varied to optimize run-times such that $N_{S}=500,50,5$ and $N_{R}=4.8\times 10^{3},4.8\times 10^{4},4.8\times 10^{5}$ for $d=1,2,3$ respectively. For all networks and dimensions tested, sampling errors were $<10^{-3}$.
From Fig.(2) it is clear that positive-P run-times scale nearly linearly with mode number for all dimensions. As the dimension increases, run-times increase by approximately an order of magnitude. For example, the largest simulated network run-times (in seconds) are $51.34$, $172.3$, and $8800$ for $d=1,2,\>\text{and}\>3$, respectively. This is unexpected, since the floating-point operations needed for the transfer matrix multiplication scale as $M^{2}$, but it indicates that memory access is more important for this range of network sizes.
Figure 2: Run-times in seconds versus mode number of positive-P simulated GCPs of dimension $d=1$ (blue), $d=2$ (orange), and $d=3$ (black). GBS networks with uniform squeezing parameter $\boldsymbol{r}=[1,\dots,1]$ and loss rate $t=0.6$ (i.e. the mean of the parameters used in Fig.(1)) are simulated for mode numbers $M=30,60,\dots,480$ where the network matrix is a Haar random unitary. The algorithm is optimised by only simulating counts within $m_{j}^{(\text{min})}\leq m_{j}\leq m_{j}^{(\text{max})}$. The set of maximum counts for $d=1,2,3$ are $\{62,88,134,218,369\}$, $\{37,50,80,115,190\}$, and $\{25,33,49,77,126\}$ respectively, while the set of minimum counts are $\{0,5,12,45,134\}$, $\{0,0,0,15,60\}$, and $\{0,0,2,11,41\}$. All positive-P simulations are performed using $E_{S}=2.4\times 10^{6}$. Sub-ensemble sizes were varied as $N_{S}=500,50,5$ and $N_{R}=4.8\times 10^{3},4.8\times 10^{4},4.8\times 10^{5}$ for $d=1,2,3$ respectively, in order to optimize run-times on the utilized hardware.
The sharp jump in time from $d=2$ to $d=3$ is due to increased memory requirements. The available memory needed to store a single complex multi-dimensional array to double-precision scales as $16N_{S}(M/d)^{d}$ bytes. Hence when $M=480$, each $d=2$ array is $\approx 0.921N_{S}$ MB, while for $d=3$ it is $\approx 65.5N_{S}$ MB. Naturally, additional memory is necessary to actually perform the computations.
Although more sensitive tests of quantum advantage are possible by comparing $d>1$ GCPs of theory and experiment, there are computational costs. Such costs are not restricted to phase-space representations. Hardware limitations are present in all quantum computing validation schemes. One must therefore balance test sensitivity with available hardware.
V Numerical simulations of experimental networks
In this section, we use the sampled positive-P method to validate experimental data of the GBS device Borealis [5]. The experimental data corresponds to the publicly available dataset of Madsen et al [5]. Other works [46, 47] have validated independently generated data from the Borealis network. These have focused on calculating the quadrature coherence scale [46] and orbits [47] of Borealis.
In this work we focus on quantitatively validating GCPs of different dimensionality, which has not been done previously. Detailed $\chi^{2}$ and Z-statistic tests are applied to compare theory and experiment for both pure and thermalized squeezed state inputs. Due to claims of quantum advantage, most of the analysis in this section is on the count patterns from the $216$-mode high squeezing (HS) parameter and $288$-mode experimental data sets. Summary tables of $\chi^{2}$ and Z-statistic test results for the Borealis data sets can be found in Appendix A.
V.1 Statistical tests
A number of statistical tests have been proposed to validate experimental GBS networks. These include total variation distance [19, 5], cross-entropy measures [19, 20, 34], two and higher-order correlation functions [48, 4], and Bayesian hypothesis testing [4, 40]. Although these tests provide useful insights into the behavior of specific correlations, they are limited to either small size systems or patterns with small numbers of detected photons, as they require computing the probabilities of specific count patterns, which is a #P-hard task.
If these patterns are closer to the ground truth distribution than classical algorithms replicating photo-detection measurements, then its assumed patterns from larger sized systems claiming quantum advantage will also pass these tests. The ground truth distribution is the target distribution GBS aims to sample from when losses are included in the transmission matrix. Generally its assumed to be close to the ideal lossless distribution, although this may not necessarily be the case.
In this work, the primary statistical test used to validate experimental data are chi-square tests [49], which in terms of GCPs are defined as [9]
$\chi^{2}=\sum_{i=1}^{k}\frac{\left(\mathcal{G}_{E,i}-\bar{\mathcal{G}}_{S,i}\right)^{2}}{\sigma_{i}^{2}}.$ (27)
Here we use the shorthand notation $\mathcal{\bar{G}}_{i}=\left\langle\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right\rangle_{P}=E_{S}^{-1}\sum_{e=1}^{E_{S}}\left(\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right)_{(e)}$ for the stochastic trajectory $e\in E_{S}$ to denote the positive-P GCP ensemble means of the $i=1,2,\dots,k$ photon count bin for the $j=1,\dots,d$ grouped count, which converges to the true theoretical GCP in the limit $\mathcal{G}_{S,i}=\lim_{E_{S}\rightarrow\infty}\bar{\mathcal{G}}_{S,i}$ for any general input state $S$. Simulated ensemble means are then compared to the experimental GCPs $\mathcal{G}_{E,i}=N_{E}^{-1}\sum_{i}c_{i}$ formed from $N_{E}\subset\mathcal{S}_{P}$ photon count patterns.
The $\chi^{2}$ tests are used for two reasons: the first is that this is a standard test applied to data from random number generators [16], where the summation is performed over $k$ count bins satisfying $\sum_{i\in S_{j}}c_{i}>10$. The second is that one can include dependence on both theoretical and experimental sampling errors as $\sigma_{i}^{2}=\sigma_{E,i}^{2}+\sigma_{T,i}^{2}$, where $\sigma_{E,i}\propto 1/\sqrt{N_{E}}$ is the estimated experimental sampling error. This provides a method of accurately determining whether observed differences are caused by large sampling error, as is the case for unbinned sparse data.
Output $\chi^{2}$ values follow a distribution that is positively skewed for small $k$. Performing a Wilson-Hilferty transformation [50], which transforms the result as $\chi^{2}\rightarrow\left(\chi^{2}/k\right)^{1/3}$, the resulting distribution converges to a Gaussian distribution $\mathcal{N}\left(\mu_{\mathcal{N}},\sigma_{\mathcal{N}}^{2}\right)$ with mean $\mu_{\mathcal{N}}=(9k-2)/9k$ and variance $\sigma_{\mathcal{N}}^{2}=2/(9k)$ for $k\geq 10$ due to the central limit theorem [50, 51]. This allows one to convert $\chi^{2}$ results into measurable probability distances using approximate Z-statistic, or Z-score, tests [52], where
$Z=\frac{\left(\chi^{2}/k\right)^{1/3}-\mu_{\mathcal{N}}}{\sigma_{\mathcal{N}}}.$ (28)
This measures how many standard deviations away a test statistic is from its expected normally distributed mean. The Z-statistic therefore determines the probability of obtaining the observed result. Agreement within sampling errors produces test results of $\chi^{2}/k\approx 1$ and $\left|Z\right|\leq 1$ when accounting for stochastic fluctuations in the output statistics. For Z-statistic tests we use a regime of validity with the limit $Z\leq 6$ to define accurate comparisons. Above this limit test statistics differ by more than $6\sigma_{\mathcal{N}}$ from their predicted means, corresponding to very small observable probabilities.
The results presented in subsequent subsections use the notation introduced previously [17]. Comparisons between theory and experiment for simulations with pure squeezed states input into the experimental $M\times M$ $\boldsymbol{T}$-matrix are denoted by the subscript $EI$, which we call the ideal ground truth distribution. Our ideal ground truth corresponds to the ground truth distributions used in Ref. [5]. This differs from the non-ideal, or modified, ground truth distribution that includes more realistic experimental effects such as thermalization, and measurement corrections in the theoretical simulations, which are denoted by the subscript $ET$.
The validation tests implemented in this paper focus on comparing experimental data with two theoretical ground truth distributions. If experimental data fail the statistical tests implemented here, it does not necessarily mean there is no possible quantum advantage present in these experiments, but it does mean that the quantum computer has generated the wrong answer relative to the target ground truth distribution.
For example, a result of $Z_{EI}\gg 1$ or $Z_{ET}\gg 1$ for GCPs of dimension $d>1$ indicates high-order correlations present in experimental data are far from their expected values for both tested ground truths. Since quantum advantage requires nonclassical correlations at high-orders, this result may indicate the tested experiment is dominated by low-order correlations, possibly increasing the likelihood of classical sampling algorithms producing samples that are closer to either distribution. The larger the binning dimension, the more sensitive the tests are to these correlations, making them harder to pass. However, at too large a dimension the experimental sampling error increases, reducing the power of the test to discriminate between different distributions.
To determine whether or not there is any quantum advantage requires comparisons with classical sampling algorithms that have a comparable or smaller error. Previous work [17] has performed some of these validation tests for a single sampling algorithm compared to threshold detector experiments. We save a comparable analysis for the PNR experiments considered in this paper for future work.
V.2 $d=1$ grouped count comparisons
First we compare the theoretical and experimental one-dimensional GCPs of the ideal ground truth distribution. Results of phase-space simulations are presented in Fig.(3a) for the $216$-mode HS data set and Fig.(4a) for the $288$-mode data set. In both cases, the simulated GCP ideal ground truth distributions appear to visually resemble the experimental data.
Visual comparisons of these distributions were also presented in [5]. When comparing our theoretical total count distributions to those obtained in [5], we found that for the $288$-mode data set, the two theoretical distributions differed slightly. After further checks by the original authors [53], the theoretical total count distribution in Fig.(4a) was determined to be the correct ideal ground truth distribution.
Despite the apparent visual agreement between theory and experiment, a quantitative $\chi^{2}$ test shows significant differences between the two distributions. The $216$-mode HS experimental data gives a result of $\chi_{EI}^{2}/k\approx 54$ with $Z_{EI}\approx 70$ Count patterns from the $288$-mode experiment have $\chi_{EI}^{2}/k\approx 411$ with $Z_{EI}\approx 167$. Although both experimental GCPs are far from their pure squeezed state distributions, the $288$-mode distribution is $98\sigma_{\mathcal{N}}$ further from its expected normally distributed mean. These numerical differences are also reflected graphically by plotting the normalized difference between distributions $\Delta\mathcal{G}_{S}^{(M)}(m)/\sigma_{i}=\left(\mathcal{G}_{E,i}-\bar{\mathcal{G}}_{I,i}\right)/\sigma_{i}$, which is plotted in Fig.(3b) for the $216$-mode HS data set and Fig.(4b) for the $288$-mode data set.
Figure 3: Comparisons of one-dimensional GCPs $\mathcal{G}_{\{1,\dots,M\}}^{(M)}(m)$ versus grouped counts $m$ for experimental data obtained from the $216$-mode HS Borealis experiment. a) Data is compared to positive-P simulations with pure squeezed state inputs, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$, and the corresponding normalised difference between distributions is plotted in b). In c) experimental patterns are now compared to the simulated GCP using thermalised squeezed state inputs with $\epsilon=0.0510\pm 0.0005$. A transmission matrix measurement correction is also applied as $t\boldsymbol{T}$, where $t=0.9941\pm 0.0005$, to improve the fit. d) The normalised difference between thermalized and experimental GCPs, which display noticeable improvement when compared to pure squeezed state simulations.
Once a thermal component $\epsilon$ is added to the input states and a transmission matrix measurement correction $t$ is applied as $t\boldsymbol{T}$, the positive-P simulated GCPs agree with the experimental distributions within sampling error. The $216$-mode HS data set requires $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ to output $\chi_{ET}^{2}/k\approx 0.88\pm 0.02$ and $\left|Z_{ET}\right|\approx 1\pm 0.2$, while the best-fit theoretical GCP for the $288$-mode data uses a similar thermal fraction of $\epsilon=0.0547\pm 0.0005$ and measurement correction $t=0.9848\pm 0.0005$ to produce $\chi_{ET}^{2}/k\approx 0.87\pm 0.02$ and $Z_{ET}\approx\left|1.1\right|\pm 0.2$. This improved agreement between theory and experiment is reflected in the normalized difference plots as seen in Fig.(3d) and Fig.(4d) which have noticeably decreased from their pure squeezed state comparisons.
Similar results are obtained for the $72$-mode and $216$-mode low squeezing (LS) experiments. In these cases, comparisons with the ideal ground truth, which output $Z_{EI}=72$ and $Z_{EI}=66$ respectively, converge to a within sampling error agreement with a thermalized theory such that $\left|Z_{ET}\right|\approx 0.2\pm 0.2$ for the $72$-mode experiment and $\left|Z_{ET}\right|\approx 0.89\pm 0.1$ for the $216$-mode LS experiment.
Figure 4: Description follows Fig.(3) except experimental count patterns are from the $288$-mode Borealis network. Input thermalised squeezed state simulations are performed using a decoherence component of $\epsilon=0.0547\pm 0.0005$ and matrix measurement error $t=0.9848\pm 0.0005$.
V.3 $d=2$ grouped count comparisons
To determine whether this excellent agreement between theoretical and experimental GCPs once decoherence and measurement corrections are accounted for continues, we increase the binning dimension to $d=2$. In this case, the $216$-mode HS experiment, whose ideal and thermalized $d=2$ distributions are shown in Fig.(5), outputs statistical test results of $Z_{EI}=58$ and $Z_{ET}=9\pm 1$.
Although the $Z_{ET}$ results are $\approx 8\sigma_{\mathcal{N}}$ larger than their $d=1$ counterparts, the ideal ground truth comparisons are $\approx 12\sigma_{\mathcal{N}}$ closer to the two-dimensional distribution than the total count comparisons, presumably because of the smaller number of counts per bin. For the $72$-mode, $216$-mode LS and $288$-mode experiments, this isn’t necessarily the case (see Appendix A), however all are closer to the thermalized ground truth than was observed in previous work [17] with threshold detector experiments.
Figure 5: Surface plots of two-dimensional GCP comparisons of theory and experiment for the $216$-mode HS experimental count patterns. Figures in the left column (a and c) plot of the full bivariate distribution $\mathcal{G}_{\{1,\dots,108\},\{109,\dots,216\}}^{(M)}(m_{1},m_{2})$, while the figures in the right column (b and d) plot the normalised difference $\Delta\mathcal{G}_{\{1,\dots,108\},\{109,\dots,216\}}^{(M)}(m_{1},m_{2})/\sigma$. In a) and b) the theoretical ideal ground truth is simulated, and in c) and d) the theoretical distribution is the thermalized ground truth. Thermalization component $\epsilon$ and measurement correction $t$, as well as the positive-P sample size are the same as in Fig.(3).
In these threshold detector experiments, $N_{E}\approx 4\rightarrow 5\times 10^{7}$ count patterns were generated and it was found that once the experimental data was binned in $d=4$ dimensions, there were too few counts per bin. This causes the bins to become more sparse, leading to a dramatic increase in experimental sampling errors that dominate statistical testing, outputting artificially small results. Therefore for these threshold detector experiments, reliable validation tests could only be performed for $d<4$ [17].
Almost all data sets from Borealis generated only $N_{E}\approx 1\times 10^{6}$ patterns. In this case, sampling errors dominate statistical testing at $d=2$, reducing the reliability of these tests such that only the $d=1$ comparisons are reliable.
The exception to this is the $16$-mode experiment whose $N_{E}\approx 8.4\times 10^{7}$ samples were used benchmark the system architecture through total variation distance tests. The network is small enough that count pattern probabilities can be computed up to a total of six detected photons. As the computational cost of computing probabilities grows with the total photon number, the full photon distribution was not computed in the analysis of Ref.[5].
Interestingly, the one-dimensional GCP for this data set is further from its ideal ground truth distribution than most experiments performed using the Borealis quantum network, where $\chi_{EI}^{2}/k\approx 2.5\times 10^{3}$ and $Z_{EI}=163$. Unlike data sets claiming quantum advantage, including decoherence and $\boldsymbol{T}$-matrix corrections is no longer enough to obtain agreement within sampling error between theoretical and experimental GCPs, instead producing statistical test results of $\chi_{ET}^{2}/k\approx 14\pm 2$ and $Z_{ET}\approx 18\pm 1$, although the difference here is mostly from the increased sample numbers.
An increase in the GCP dimension to $d=2$ shows the $16$-mode data is even further from both the ideal and non-ideal ground truth distributions as can be seen in Fig.(6), outputting $Z_{EI}=301$ and $Z_{ET}=86\pm 3$. Additional $d=2$ comparisons can be performed by randomly permuting the count patterns as each permutation allows one to compare different binned correlation moments. As some correlation moments may be closer to their theoretical values than others, Z-scores are likely to vary between permutations, as was observed in threshold detector experiment comparisons [17]. Focusing on comparisons with the non-ideal ground truth, a mean increase in the resulting Z-scores is observed from ten permutations, such that $\left\langle Z_{ET}\right\rangle_{\mathcal{RP}}\approx 139$, where $\left\langle\dots\right\rangle_{\mathcal{RP}}$ denotes a mean over random permutations. Although some permutations are closer to the non-ideal ground truth, with the closest being $Z_{ET}\approx 73$, the majority are further, where the largest observed Z-score was $Z_{ET}\approx 235$ from a single permutation.
Although this is not unexpected, such results indicate further systematic errors besides those tested are present. These may include unequal losses in each network channel, phase noise, shot-to-shot fluctuations, miscounted photons, or detector efficiency fluctuations, among others. We conjecture that the presence of further systematic errors other than decoherence and measurement errors are also present in the $216$-mode HS and $288$-mode implementations. However the small sample size means such errors cannot be resolved in statistical testing for these data sets, limiting a thorough analysis.
Figure 6: Description follows Fig.(5) however in this case experimental count patterns are from the $16$-mode Borealis network, such that the $d=2$ GCP is $\mathcal{G}_{\{1,\dots,8\},\{9,\dots,16\}}^{(M)}(m_{1},m_{2})$. For this network, the thermalization component is $\epsilon=0.0648\pm 0.0005$ while the matrix measurement error is $t=0.9841\pm 0.0005$. All simulations used $E_{S}=1.2\times 10^{6}$ samples.
V.4 Photon number moments
Finally, for completeness, we simulate the mean output photon number moments $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ for all $j$ modes, where summary tables of statistical test results for all data sets is also presented in Appendix A. As with the GCP comparisons, count patterns from the $216$-mode HS and $288$-mode experiments are far from their theoretical first-order moments for pure squeezed state inputs with $Z_{EI}\approx 86$ and $Z_{EI}\approx 192$ for the $216$-mode HS and $288$-mode data sets respectively.
Unlike the GCP comparisons, decoherence and transmission matrix corrections no longer cause the theoretical moments to be within the sampling error-bars of the experimental moments, where $Z_{ET}\approx 45$ for the $216$-mode HS data set and the $288$-mode data set is $Z_{ET}\sigma_{\mathcal{N}}\approx 63\sigma_{\mathcal{N}}$ standard deviations away from its expected mean. Similar trends are present for all data sets where, as was the case for GCPs, the $16$-mode moments are furthest from their expected first order moments than any other experiment. We emphasize that statistical testing is more accurate for this data set due to the large number of experimental samples, allowing one to resolve finer differences between theoretical and experimental photon number moments.
One noticeable feature in the experimental and theoretical moments is the decrease in $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ as $j\rightarrow M$, with sharp decreases occurring for later modes as seen in Fig.(7) for $216175<j<$ 216 and Fig.(8) for $288250<j<$ 288, while distinct sharp spikes in the moments occur every $16$ modes. This clear structure is present in all data sets from this experiment but is not present in previous implementations with threshold detectors [3, 4] and is likely due to the time-domain multiplexing scheme used to form the linear network in Ref. [5]. The experimental set-up interferes a train of input squeezed photons on multiple fibre delay lines separated by temporal bins of various dimension. Output pulses are then sent into a 1-to-16 mode demultiplexer before measurements are performed by PNR detectors.
Figure 7: Upper panel: Input thermalised squeezed state theoretical mean output photon numbers $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ versus mode number $j$ are compared to their experimental counterparts for data from the $216$-mode HS Borealis experiment, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$. Decoherence and $\boldsymbol{T}$-matrix parameters are $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ which, although improving the fit with experimental data, no longer cause theory and experiment to agree within sampling error. Lower panel: Corresponding normalised difference between marginal moments.
Considering its input-output ratio the demultiplexer may be the cause of the periodic spikes. However the structure in the form of the overall decrease in $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ for increasing mode number may be due to the use of fibre delay lines. Although providing a more compact architecture, it comes at the expense of reduced connectivity, i.e. photon interference between all modes isn’t possible. This has the effect of introducing a structure to the output photons that can be exploited by classical samplers such as the tree-width sampler [54].
Figure 8: Figure description follows Fig.(7), except data is now from the $288$-mode Borealis experiment where corresponding fitting parameters are $\epsilon=0.0547\pm 0.0005$ and $t=0.9848\pm 0.0005$.
VI Conclusion
In this paper, we extended a binning method developed by Mandel [11, 12] to simulate nonclassical photon statistics of photon-number resolving detectors using the positive P-representation. We performed validation tests on quantum computers claiming quantum advantage with photon-number resolving detectors.
This method is computationally efficient, with computational run-times of a few minutes on a 12-core desktop at $\sim 10^{11}$ floating point operations/ sec (Linpack Flops) for one-dimensional GCP simulations of the largest sized networks implemented in the paper. Since the positive P-representation accurately computes correlation moments to any order generated in the photonic network, it is a time and resource efficient method of validation compared to classical algorithms that compute the Hafnian to replicate photo-detection measurements, or tensor network approximations [32].
Increasing the GCP dimension causes an increase in computation time, although this is still significantly faster than other validation methods. Directly calculating $100\times 100$ Hafnians are estimated [32] to take $14$ hours on Fugaku, one of the worlds largest supercomputers in 2024, at $\sim 10^{18}$ Flops. This would compute one probability, but rounding errors are unknown and are likely to be extremely large at current $64$ bit numerical precision. Generating enough classical counts for validation using this approach would take $\sim 10^{16}s$ [32] on Fugaku, indicating an overall speed-up of $>10^{18}$ using the positive-P method at these network sizes.
Our method does not generate classical counts, so it is strictly for validation of the computational task, and does not replicate the required random counts. Validation tests were performed on GBS data from a recent experiment claiming quantum advantage [5]. We showed that within sampling error agreement between theory and experiment for one-dimensional, but not two-dimensional GCPs is obtained once decoherence is included in the input states, and a transmission matrix measurement error correction is applied. There are large discrepancies from $\chi^{2}$ tests on the individual channel count probabilities, because these can be calculated and measured very accurately. This result agrees with a trend obtained previously for threshold detector experiments [9, 17].
Comparisons with threshold detector experiments are slightly misleading, however. Due to the smaller number of experimental samples from data sets claiming quantum advantage using PNR methods, only a small fraction of the possible tests could be performed. This was because experimental sampling errors were too large to apply more sensitive tests. Based on comparisons of the $16$-mode experimental data, whose sample size was large enough to perform accurate statistical testing, we conjecture that more experimental errors are present in the quantum advantage data sets than those tested for in this paper.
In the longer term, due to its computational efficiency and ability to generate any measurable moment, the positive-P based binning method could be used as a means to provide error correction, either by correcting parameter values as shown in this paper, or by giving feedback so they can be adjusted. We suggest that this could provide a more practical route to quantum computational advantage than only using improved hardware, as is also thought to be the case for logic gate based quantum computers.
Finally, even though we focused here on GBS networks with pure or thermalized squeezed states, both the binning method and positive P-representation can be used to simulate photon statistics of any nonclassical state.
Acknowledgments
ASD would like to thank Nicholas Quesada, Fabian Laudenbach and Jonathan Lavoie for helpful discussions. This research was funded through grants from NTT Phi Laboratories and a Templeton Foundation grant ID 62843.
Appendix A: Summary of all experimental comparisons
The Borealis GBS implemented in [5] performed experiments for varying system sizes of $M=16,72,216,\>\text{and}\>288$. The $216$-mode experiment was repeated for the mean squeezing parameters $\bar{r}=1.103$ and $\bar{r}=0.533$, hence they are distinguished as high squeezing (HS) and low squeezing (LS) experiments, respectively. The ideal ground truth distributions for each data set are formulated from their corresponding lossy transmission matrices $\boldsymbol{T}$ and simulated pure squeezed states using squeezing vectors $\boldsymbol{r}$.
The graphs presented in Section V of the main text compare theory and experiment predominantly for the $216$-mode HS and $288$-mode data sets, due to claims of quantum advantage in the original paper. Here we summarize statistical test results for all data sets in [5], starting with one-dimensional GCPs for the ideal ground truth in Table.(1). We use the subscripts $EI$ and $ET$ to denote comparisons between experimental data and the ideal and non-ideal ground truth distributions, respectively. These results clearly show that there are very large discrepancies in all data-sets, and that these differences cannot be explained by the intrinsic sampling error from using a finite number of samples.
A summary of the non-ideal ground truth comparisons are presented in Table (2), where the $t$ and $\epsilon$ fitting parameters are found using a Nelder-Mead simplex algorithm which minimizes the distance between experimental and theoretical distributions. As stated earlier, experimental data agrees with our modified theoretical predictions within sampling error for all data sets bar the $16$-mode experiment.
Mode number $\chi_{EI}^{2}/k$ $k$ $Z_{EI}$
$16$ $2.53\times 10^{3}$ $37$ $163$
$72$ $171$ $56$ $72$
$216$-(LS) $165$ $48$ $66$
$216$-(HS) $54$ $140$ $70$
$288$ $411$ $149$ $167$
Table 1: Summary of statistical tests for comparisons of one-dimensional GCPs $\mathcal{G}_{\{1,\dots,M\}}^{(M)}(m)$ for all experimental data sets obtained from [5]. Chi-square and Z-statistic tests are generated from positive-P simulated GCPs with an ensemble size of $E_{S}=1.2\times 10^{6}$. Although the $16$-mode and $288$-mode data are furthest from the ideal ground truth distribution, all experimental count patterns are at least $66\sigma_{\mathcal{N}}$ from their expected normally distributed means.
Mode number $t$ $\epsilon$ $\chi_{ET}^{2}/k$ $k$ $\left|Z_{ET}\right|$
$16$ $0.9841$ $0.0648$ $14\pm 2$ $36$ $18\pm 1$
$72$ $0.9861$ $0.0507$ $1.02\pm 0.03$ $54$ $0.2\pm 0.2$
$216$-(LS) $0.9882$ $0.0105$ $0.82\pm 0.02$ $47$ $0.89\pm 0.1$
$216$-(HS) $0.9941$ $0.0510$ $0.88\pm 0.02$ $371$±0.2plus-or-minus10.21$\pm$ 0.21 ± 0.2
$288$ $0.9848$ $0.0547$ $0.87\pm 0.02$ $142$ $1.1\pm 0.2$
Table 2: Description follows those presented in Table.(1) except now experimental data is compared to a modified non-ideal ground truth distribution for the one-dimensional GCPs. Thermalisation component $\epsilon$ and transmission matrix measurement correction $t$ are used to modify the ground truth distribution to best fit the experimental data, where $\epsilon,t$ each have error bars of ±$\pm 0.0005$ for all data sets. The $\chi_{ET}^{2}/k$ and $Z_{ET}$ values are averages over ten simulation runs with corresponding standard deviations.
Comparisons of the two-dimensional GCPs and mean output photon number moments are presented in Table.(Appendix A: Summary of all experimental comparisons) and Table.(4), respectively, for the ideal and non-ideal ground truth distributions. These show large differences in the two-dimensional GCPs compared to the ideal ground-truth. However, the two-parameter fitting model used in the one-dimensional comparisons is insufficient to give agreement in the thermalized case.
Mode number $Z_{EI}$ $k$ $\left|Z_{ET}\right|$ $k$
$16$ $301$ $426$ $86\pm 3$ $416$
          $72$ $82$ $772$ $12\pm 1$ $736$
          $216$-(LS) $75$ $658$ $10\pm 1$ $640$
          $216$-(HS) $58$ $4326$ $9\pm 1$ $4154$
          $288$ $203$ $4877$ $11\pm 1$ $4571$
Table 3: Summary of statistical tests for comparisons of two-dimensional GCPs $\mathcal{G}_{\{1,\dots,M/2\},\{M/2+1,\dots,M\}}^{(M)}(m_{1},m_{2})$ for all experimental data sets obtained from [5]. Chi-square and Z-statistic tests are generated from positive-P simulated GCPs with an ensemble size of $E_{S}=1.2\times 10^{6}$. The $\epsilon$ and $t$ fitting parameters are the same those used in Table.(2), where $Z_{ET}$ values are averages over ten simulation runs with corresponding standard deviations.
Mode number $Z_{EI}$ $Z_{ET}$
$16$ $264$ $243$
$72$ $78$ $17$
$216$-(LS) $85$ $37$
$216$-(HS) $86$ $45$
$288$ $192$ $63$
Table 4: Summary of Z-statistic test results for comparisons of both the ideal and non-ideal ground truth distributions for the mean output photon number $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ of each experimental data set. Positive-P simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$ for all data sets except the $16$-mode experiment, where $E_{S}=3.36\times 10^{7}$ is needed in order for $\sigma_{T,j}\apprle\sigma_{E,j}$ which is required for accurate simulations. Fitting parameters applied to each data set are obtained from Table.(2).
The last table, giving the comparisons for the mean photon number per channel, has a much lower sampling error than the other results tabulated, since all the output samples can be used to obtain these mean values. This shows discrepancies outside the sampling error limits, indicating that transmission matrix parameter errors could be a significant part of the overall differences. This is consistent with the increased errors found for the two-dimensional GCPs.
Appendix B: Estimating sampling errors of moments
Accurately estimating sampling errors is essential when comparing theoretical simulations with experimental data obtained via sampling. This is particularly important when using statistical tests that are dependent on sampling errors, as incorrect estimates will result in misleading test results.
Theoretical sampling errors are estimated numerically by subdividing the total phase-space ensemble as $E_{S}=N_{S}N_{R}$. Breaking the full ensemble of phase-space samples into sub-ensembles is a common numerical procedure for applications of phase-space representations yielding two complementary benefits: efficient multi-core parallel computing, decreasing computation time depending on available resources, and estimates of theoretical sampling errors [43].
The first sub-ensemble corresponds to sampling from the Gaussian phase-space distribution $N_{S}$ times and $N_{R}$ is the number of times the computation is repeated. Provided $N_{R}\gg 1$, the second sub-ensemble allows one to accurately estimate the standard deviation in the ensemble mean as $\sigma_{T,i}=\sigma_{t,i}/\sqrt{N_{R}}$ [52, 55], where $\sigma_{t}$ is the computed sample standard deviation which takes the general form [43]
$\sigma_{t}=\sqrt{\frac{1}{N_{R}-1}\sum_{i=1}^{N_{R}}\left(\bar{O}^{(i)}-\bar{O}\right)}.$ (29)
Here, $O(\boldsymbol{\alpha},\boldsymbol{\beta})$ is some general phase-space observable, $\bar{O}^{(i)}=N_{S}^{-1}\sum_{k=1}^{N_{S}}(O^{(k)})^{(i)}$ denotes the first sub-ensemble mean for the $i\in N_{R}$ trajectory and $\bar{O}=E_{S}^{-1}\sum_{e=1}^{E_{S}}O_{(e)}$ is the full ensemble mean. In statistics literature [52, 56], $\sigma_{T,i}$ is commonly called the standard error of the mean and its valid for both binned distributions and moments.
For experimental sampling errors defined as $\sigma_{E,i}=\sigma_{e,i}/\sqrt{N_{E}}$, we are interested in estimating the standard deviation of the experimental sample means $\sigma_{e,i}$. Unlike theoretical sampling errors this isn’t performed numerically. Instead, the variances of GCPs for both threshold and PNR detector experiments have Poissonian fluctuations, such that $\sigma_{e,i}=\sqrt{\mathcal{G}_{S,i}}$ for some general input state $S$. Since the true theoretical GCP $\mathcal{G}_{S,i}$ is seldom known, we estimate it in phase-space as $\mathcal{G}_{S,i}=\lim_{E_{S}\rightarrow\infty}\bar{\mathcal{G}}_{S,i}$.
However, estimating $\sigma_{e,i}$ for moments (either raw or central) and cumulants is somewhat more complicated as these aren’t probabilities but instead tell one about various properties of a probability distribution. In this appendix, we describe the method of estimating experimental sampling errors of raw moments used in the comparisons presented in the main text. Details on how to estimate sampling errors for central moments and cumulants can be found in [56, 57, 58].
Sampling errors of raw moments
Let $Y$ be an i.i.d random variable with probability density function $P(y)$, then the $n$-th raw moment is defined as
$\mu^{\prime}_{n}=\left\langle Y^{n}\right\rangle=\int y^{n}P(y)\text{d}y.$ (30)
Suppose one generates $\mathcal{S}$ samples denoted $y_{1},\dots,y_{\mathcal{S}}$ from the distribution $P(y)$. The $n$-th sample raw moment is then defined as [56, 57, 58]
$m^{\prime}_{n}=\frac{1}{\mathcal{S}}\sum_{i=1}^{\mathcal{S}}y_{i}^{n},$ (31)
the expected value of which is an unbiased estimator of $\mu^{\prime}_{n}$
$\left\langle m^{\prime}_{n}\right\rangle=\frac{1}{\mathcal{S}}\sum_{i=1}^{\mathcal{S}}\left\langle y_{i}^{n}\right\rangle=\mu^{\prime}_{n}.$ (32)
For our purposes, we are interested in the variance $\sigma_{m^{\prime}_{n}}^{2}=\left\langle(m^{\prime}_{n})^{2}\right\rangle-\left\langle m^{\prime}_{n}\right\rangle^{2}$. Following from Ref. [56], the expectation value of the squared moment is simplified as
$\left\langle(m^{\prime}_{n})^{2}\right\rangle$ $=\frac{1}{\mathcal{S}^{2}}\left\langle\sum_{i=1}^{\mathcal{S}}\sum_{j=1}^{\mathcal{S}}y_{i}^{n}y_{j}^{n}\right\rangle$
$=\frac{1}{\mathcal{S}^{2}}\left\langle\sum_{l=1}^{\mathcal{S}}y_{l}^{2n}+\sum_{i\neq j}^{\mathcal{S}}y_{i}^{n}y_{j}^{n}\right\rangle,$ (33)
where the summations have been expanded to include instances of equal $l=i=j$ and non-equal $i\neq j$ samples. Utilizing the simplifications $\sum_{i=1}^{\mathcal{S}}\left\langle y_{i}^{n}\right\rangle=\mathcal{S}\mu^{\prime}_{n}$ from Eq.(32) and $\sum_{i\neq j}^{\mathcal{S}}=\sum_{i}^{\mathcal{S}-1}\sum_{j}^{\mathcal{S}}$, the second-order sample raw moment is simplified as
$\left\langle(m^{\prime}_{n})^{2}\right\rangle=\frac{\mu^{\prime}_{2n}+(\mathcal{S}-1)(\mu^{\prime}_{n})^{2}}{\mathcal{S}},$ (34)
allowing one to estimate the variance in the sample raw moment as [56]
$\sigma_{m^{\prime}_{n}}^{2}=\frac{\mu^{\prime}_{2n}-(\mu^{\prime}_{n})^{2}}{\mathcal{S}}.$ (35)
Its possible to extend the above description for univariate statistics to the multivariate case. Let $\boldsymbol{Y}=[Y_{1},Y_{2},\dots]$ be a vector of i.i.d random variables with multivariate probability density $P(\boldsymbol{y})$, where $\boldsymbol{y}=[y_{1},y_{2},\dots]$. The $n$-th raw moment is then defined as
$\mu^{\prime}_{n_{1},n_{2},\dots}=\left\langle Y_{1}^{n_{1}}Y_{2}^{n_{2}}\dots\right\rangle=\int\left(y_{1}^{n_{1}}y_{2}^{n_{2}}\dots\right)P(\boldsymbol{y})\text{d}\boldsymbol{y}.$ (36)
As in the univariate case, suppose one generates $\mathcal{S}$ samples of the $j=1,2,\dots$-th random variable in $\boldsymbol{Y}$, denoted $y_{1,j},\dots,y_{\mathcal{S},j}$, then the multivariate $n$-th sample raw moment is defined as [57]
$m^{\prime}_{n_{1},n_{2},\dots}=\frac{1}{\mathcal{S}}\sum_{i=1}^{\mathcal{S}}y_{i,1}^{n_{1}}y_{i,2}^{n_{2}}\cdots,$ (37)
whose expected value is an unbiased estimate of the multivariate raw moment [58]
$\left\langle m^{\prime}_{n_{1},n_{2},\dots}\right\rangle=\frac{1}{\mathcal{S}}\sum_{i=1}^{\mathcal{S}}\left\langle y_{i,1}^{n_{1}}y_{i,2}^{n_{2}}\cdots\right\rangle=\mu^{\prime}_{n_{1},n_{2},\dots}.$ (38)
Following from the derivation of the univariate case, the variance of the multivariate raw sample moment is obtained as [59]
$\sigma_{m^{\prime}_{n_{1},n_{2},\dots}}^{2}=\frac{\mu^{\prime}_{2n_{1},2n_{2},\dots}-(\mu^{\prime}_{n_{1},n_{2},\dots})^{2}}{\mathcal{S}}.$ (39)
By setting $n_{1}=n_{2}=\dots=1$, sampling errors for higher photon number moments such as $\left\langle\hat{n}^{\prime}_{j}\hat{n}^{\prime}_{k}\right\rangle$, $\left\langle\hat{n}^{\prime}_{j}\hat{n}^{\prime}_{k}\hat{n}^{\prime}_{l}\right\rangle$ can now be estimated.
Accuracy of sampling error estimates
In this paper, the raw moments of interest are the mean output photon numbers per mode $\left\langle\hat{n}^{\prime}_{j}\right\rangle$. Using Eq.(35), the experimental sampling errors can therefore be estimated as
$\sigma_{E,j}=\frac{\sigma_{e,j}}{\sqrt{N_{E}}}=\sqrt{\frac{\left\langle\hat{n}_{j}^{\prime 2}\right\rangle-\left\langle\hat{n}^{\prime}_{j}\right\rangle^{2}}{N_{E}}},$ (40)
where, as was the case for GCPs, we estimate the true theoretical photon number moments in phase-space as $\left\langle\hat{n}^{\prime}_{j}\right\rangle=\lim_{E_{S}\rightarrow\infty}\bar{n}^{\prime}_{j}$. To determine the accuracy of this estimate, we must generate photon count patterns corresponding to PNR detectors from an exactly known model and compare the estimated sampling error with the standard deviation of the mean output photon number of the count patterns from their exact values. To do this, the simplest method is to use classical states.
Classical state PNR photo-detection algorithm
In the case of $N$ classical states input into a GBS linear network, one has rotated to a classical phase-space and we can use the coherent amplitudes corresponding to the diagonal P-representation to replicate photo-detection measurements of PNR detectors.
Generally, each stochastic trajectory in $E_{S}$ corresponds to a single experimental shot. Therefore one can conserve the classical correlations generated within the network by computing the phase-space observable projector of the $j$-th detector $p_{j}$, which corresponds to the operator Eq.(5), for the $k$-th trajectory as
$(p_{j}(c_{j}^{f}))^{(k)}=\frac{1}{c_{j}^{f}!}\left(\left(n^{\prime}_{j}\right)^{c_{j}^{f}}e^{-n^{\prime}_{j}}\right)^{(k)},$ (41)
where $k\in E_{S}$, $n^{\prime}_{j}=\left|\alpha_{j}\right|^{2}$ in the diagonal P-representation and $c_{j}^{f}=0,1,2,\dots,c_{j}^{(\text{max}),f}$ is a photon count of our fake experiment up to some maximum cut-off $c_{j}^{(\text{max}),f}$.
The phase-space projectors determine the probability of observing a specific count $c_{j}^{f}$. Since each detector can distinguish between multiple oncoming photons, the actual computation involves calculating a vector of probabilities for each stochastic trajectory:
$\boldsymbol{P}_{j}^{(k)}=[P_{0,j}^{(k)},P_{1,j}^{(k)},\dots,P_{\text{max},j}^{(k)}],$ (42)
where $P_{0,j}^{(k)}=(p_{j}(0))^{(k)}$ is the success probability of observing no counts at the $j$-th detector, $P_{1,j}^{(k)}=(p_{j}(1))^{(k)}$ is the success probability of observing a single count at the $j$-th detector, and so on. The PNR photo-detection measurements are then replicated by randomly sampling Eq.(42), which outputs the fake photon count vector
$(\boldsymbol{c}^{f})^{(k)}=[P_{c_{1},1}^{(k)},P_{c_{2},2}^{(k)},\dots,P_{c_{M},M}^{(k)}].$
As the method scales with the total phase-space ensemble size, the number of faked patterns generated is $N_{F}=E_{S}$. These are binned in the same way as the experimental data, where $m_{j}^{f}=\sum_{i\in S_{j}}c_{i}^{f}$ is the fake grouped count corresponding to the GCP $\mathcal{G}_{j}^{f}=N_{F}^{-1}\sum_{i\in S_{j}}c_{i}^{f}$.
Tests of estimate accuracy
The photon counting distributions for a number of classical states can be exactly computed, one of which is multi-mode thermal states. The one-dimensional GCP of a multi-mode thermal state is obtained from the single-mode distribution [26]
$\mathcal{G}_{S}^{(M)}(m)=p(1-p)^{m},$ (43)
where $p=1/(1+n)$ is the success probability requiring a uniform squeezing parameter and hence equal photon numbers $n=n_{1}=\dots=n_{N}$. The single-mode theory follows a geometric distribution which is a special case of the negative binomial distribution: if $X$ is an independent and geometrically distributed random variable with parameter $p$, then $Y=\sum_{j}^{M}X_{j}$ follows a negative binomial distribution with parameters $M$ and $p$.
The one-dimensional GCP theory of multi-mode thermal state inputs therefore follows a negative binomial distribution
$\mathcal{G}_{S}^{(M)}(m)=\left(\begin{array}[]{c}m+M-1\\ m\end{array}\right)p^{M}(1-p)^{m}.$ (44)
For the case of thermal photons transformed by a Haar random unitary, the input and output mean photon numbers are equal $n=\left\langle\hat{n}^{\prime}\right\rangle=\sinh^{2}(r)$ and the photon number variance can be computed exactly as $\sigma_{n}^{2}=n\left(n+1\right)=\sinh^{2}(r)\left(\sinh^{2}(r)+1\right)$ [60].
Now that we have a method of generating photon counts and an exactly known model with which to compare those counts, we calculate the accuracy of the estimate Eq.(40) for $M=N=50$ thermal states with uniform squeezing $\boldsymbol{r}=[0.5,\dots,0.5]$ sent into a lossless linear network described by a Haar random unitary matrix $\boldsymbol{U}$. We generate $N_{F}=4\times 10^{6}$ count patterns using a detector cut-off of $c_{j}^{(\text{max}),f}=13$ and bin them to compute the mean output photon number $\bar{n}_{j}^{\prime f}$, which is compared to the exact value in Fig.(9).
Following from Eq.(40), the sampling error of our fake experiment is estimated by setting $\sigma_{e}=\sigma_{n}=\sqrt{n(n+1)}$, such that
$\sigma_{E}=\sqrt{\frac{n(n+1)}{N_{F}}}.$ (45)
If this is an accurate estimate, it should converge to the actual error between moments
$\sigma_{A}=\sqrt{\frac{1}{M}\sum_{j=1}^{M}\left(\bar{n}_{j}^{\prime f}-n\right)^{2}},$ (46)
which is simply the standard deviation as $n$ is our expected mean value.
For the data in Fig.(9), our estimated sampling error is $\sigma_{E}\approx 2.938\times 10^{-4}$ while the actual error is $\sigma_{A}\approx 3.095\times 10^{-4}$, corresponding to an $\approx 95\%$ agreement. If instead we assumed one could estimate the errors as Poissonian fluctuations, such as for GCPs, then $\sigma_{E}=\sqrt{n/N_{F}}\approx 2.605\times 10^{-4}$ which is only $\approx 84\%$ of the actual error.
Therefore, Eq.(40) is an accurate method of estimating experimental sampling errors of moments. Naturally, the larger the number of experimental samples the better the estimate will be as the experimental moments will converge to their ground truth values more readily.
Figure 9: The exact mean photon number (solid black line) of $N=50$ thermal states sent into a lossless linear network compared to the mean photon number (solid orange dots) of a classical algorithm that replicates PNR measurements of classical states using the diagonal P-representation. The $N_{F}=4\times 10^{6}$ count patterns are binned to form the observable moments, and the standard deviation of these moments from their exact values $\sigma_{A}$ is the sampling error estimated using Eq.(45).